I am stuck on this step of finding the area bounded by the following curves.

y = 16-x^2, y = 8x - x^2, x = 1, x = 3

I can figure everything out until I get to the point of combining A=Aleft = Aright (shown below)
Answer #1

= 16-0² =) A= Aleft & Aright y= , y = 8x-x² > -1, X-3 intersecting at then 16-Y?= 82-42 8x=16 -) А 3 (8x-x2) ate x=2 2. 7 16-